OpenStudy (anonymous):

Write an equation of the line given the following conditions: Perpendicular to 2x-3y=9 & passes through the point (-3,1).

OpenStudy (anonymous):

write in regular form the first equation: 2x-9=3x => y=2/3x-3 perpendicular mean that the slope is the "opponent of the inverse" so the new m is -(1/m) => y=-(3/2)x+q Now... I leave q because I dont' know what of the infinite parallel lines is the solution, to find it, just resolve the equation with the x-y of the point, this mean: 1 = -(3/2)*(-3)+q => 1=9/2+q => q=1-9/2 => q=-7/2 so finally, your line have equation of the form y=-(3/2)x-7/2 !! ;)

OpenStudy (anonymous):

*the first equation is 2x-9=3y